# Spatial Transformers for Radio Map Estimation

### v1

@@ Abstract
Radio map estimation (RME) involves spatial interpolation of radio measurements to predict metrics such as the received signal strength at locations where no measurements were collected. The most popular estimators nowadays project the measurement locations to a regular grid and complete the resulting measurement tensor with a convolutional deep neural network. Unfortunately, these approaches suffer from poor spatial resolution and require a great number of parameters. The first contribution of this paper addresses these limitations by means of an attention-based estimator named Spatial TransfOrmer for Radio Map estimation (STORM). This scheme not only outperforms the existing estimators, but also exhibits lower computational complexity, translation equivariance, rotation equivariance, and full spatial resolution. The second contribution is an extended transformer architecture that allows STORM to perform active sensing, where the next measurement location is selected based on the previous measurements. This is particularly useful for minimization of drive tests (MDT) in cellular networks, where operators request user equipment to collect measurements. Finally, STORM is extensively validated by experiments with one ray-tracing and two real datasets.
@@
Radio map estimation, transformers, attention-based learning, wireless communications. This work has been funded by the IKTPLUSS grant 311994 of the Research Council of Norway.
@@ I Introduction
Radio maps (cf. Fig. 1), also known as radio environment maps, provide radio frequency (RF) metrics such as the received signal strength across a geographical region [1, 2]. Radio maps find a large number of applications, including network planning, frequency planning, cellular communications, device-to-device communications, dynamic spectrum access, robot path planning, aerial traffic management in unmanned aerial systems, and fingerprinting localization to name a few; see e.g. [3, 4, 1, 5] and references therein.
Radio map estimation (RME) involves constructing a radio map by relying on measurements collected across the area of interest. Before the advent of deep learning, the most popular estimators were built upon kernel-based learning (see [6] and references therein), Kriging [7, 8], sparsity-based inference [9], matrix completion [10], dictionary learning [11], and graphical models [12]. The most recent estimators are based on deep neural network (DNNs); see e.g. [13, 14, 15, 8]. Unfortunately, these schemes entail grid discretization and a very large number of trainable parameters, which renders them computationally expensive and drastically limits their spatial resolution. Besides, they lack important desirable properties in RME, such as translation and rotation equivariance.
@@
Some works related to RME have considered attention-based estimators, which is the topic of this paper. For example, vision transformers have been used to accommodate side information, such as building maps [16, 17] and satellite images [18]. In other works, transformers are fed with radio measurements. This is the case of [19], which uses a transformer for predicting what a device would measure given the measurements collected by a device with different hardware characteristics, and of [20], where a transformer fills missing RSS features for fingerprinting-based localization. Thus, transformers have been applied to problems that are related to RME or to modified versions of the RME problem where transformers are used to process images. However, the plain RME problem, where a radio map needs to be constructed relying on radio measurements and their locations, requires spatial interpolation of radio measurements and this has never been tackled using transformers.
Figure 1: Example of a radio map estimate obtained with the proposed STORM estimator in the USRP dataset (cf. Sec. VI).
The first contribution of this paper is an attention-based scheme referred to as Spatial TransfOrmer for Radio Map estimation (STORM). As shown by numerical experiments in one ray-tracing and two real datasets, STORM outperforms the existing estimators, thereby setting the state of the art in RME. Besides, it offers key advantages over existing DNN estimators: First, its complexity is significantly lower. For example, the RadioUnet estimators in [14] have 6 M and 25 M parameters, whereas STORM has just 100 k parameters. Second, STORM is gridless. This implies that, unlike DNN estimators, (i) its spatial resolution is not limited, (ii) it is translation and rotation equivariant, which are desirable properties in view of Maxwell’s equations, and (iii), it need not be retrained if the grid spacing changes. Third, STORM can estimate the map only at the relevant locations, whereas DNN estimators need to compute the map at all grid locations. Fourth, STORM can accommodate measurements outside the region of interest, as opposed to DNN estimators. STROM also compares favorably to Kriging, which is the non-DNN estimator of choice [21, 22]. While Kriging requires a matrix inversion, whose complexity is cubic in the number of measurements, STORM enjoys quadratic complexity.
@@ II The RME Problem
This section presents the most prevalent formulation of RME, both in the conventional and active sensing setups. For simplicity, the exposition assumes that the signal strength is quantified using the received signal power, but it carries over readily to other metrics.
@@
The received power is measured at ${{\color[rgb]{0,0,0}N}}$ locations $\left\{{\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}\subset{\color[rgb]{0,0,0}\mathcal{R}}$ by one or multiple receivers (or sensors). The ${{\color[rgb]{0,0,0}n}}$-th measurement can be written as ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}}={{\color[rgb]{0,0,0}\gamma}}\left({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\right)+{\color[rgb]{0,0,0}\zeta}_{{{\color[rgb]{0,0,0}n}}}$, where ${\color[rgb]{0,0,0}\zeta}_{{{\color[rgb]{0,0,0}n}}}$ denotes measurement error. Given $\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}$, the RME problem is to estimate ${{\color[rgb]{0,0,0}\gamma}}({\color[rgb]{0,0,0}\bm{r}})$, ${\color[rgb]{0,0,0}\bm{r}}\in{\color[rgb]{0,0,0}\mathcal{R}}$. On the other hand, in the active sensing problem, one is given the measurements $\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}$ as well as the set $\{{\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\}_{{{\color[rgb]{0,0,0}n}}={{\color[rgb]{0,0,0}N}}+1}^{{{\color[rgb]{0,0,0}N}}+{\color[rgb]{0,0,0}M}}$ of candidate locations and selects one of these candidate locations, say ${\color[rgb]{0,0,0}\bm{r}}_{{\color[rgb]{0,0,0}m}}$, to collect the next measurement ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{\color[rgb]{0,0,0}m}}$. The goal is to choose ${\color[rgb]{0,0,0}m}$ so that a target metric of the estimation error is minimized given the measurements $\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\}_{{{\color[rgb]{0,0,0}n}}={{\color[rgb]{0,0,0}N}}+1}^{{\color[rgb]{0,0,0}M}}\cup\{({\color[rgb]{0,0,0}\bm{r}}_{{\color[rgb]{0,0,0}m}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{\color[rgb]{0,0,0}m}})\}$.
@@ III-A Attention Heads
The attention weights are determined by the relation between ${\color[rgb]{0,0,0}\bm{x}}$ and the vectors in ${\color[rgb]{0,0,0}\bm{Z}}$. The learnable function ${\color[rgb]{0,0,0}\alpha}:\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{z}}}\times\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{x}}}\rightarrow\mathbb{R}_{+}$ can be thought of as quantifying the similarity between ${\color[rgb]{0,0,0}\bm{z}}$ and ${\color[rgb]{0,0,0}\bm{x}}$. In this way, if ${\color[rgb]{0,0,0}\bm{x}}$ is very similar to ${\color[rgb]{0,0,0}\bm{z}}_{{{\color[rgb]{0,0,0}n}}_{0}}$ for some ${{\color[rgb]{0,0,0}n}}_{0}$ and dissimilar to the remaining reference vectors, then ${\color[rgb]{0,0,0}\alpha}({\color[rgb]{0,0,0}\bm{z}}_{{{\color[rgb]{0,0,0}n}}_{0}},{\color[rgb]{0,0,0}\bm{x}})$ will dominate and ${\color[rgb]{0,0,0}H}({\color[rgb]{0,0,0}\bm{Z}},{\color[rgb]{0,0,0}\bm{x}})\approx{\color[rgb]{0,0,0}v}({\color[rgb]{0,0,0}\bm{z}}_{{{\color[rgb]{0,0,0}n}}_{0}})$. Usually, ${\color[rgb]{0,0,0}\alpha}$ is the so-called inner-product attention function [^1] A factor of $1/\sqrt{{\color[rgb]{0,0,0}D}}$ is often explicitly included inside the exponential but it is absorbed here into either ${\color[rgb]{0,0,0}k}$ or ${\color[rgb]{0,0,0}q}$ to simplify notation. ${\color[rgb]{0,0,0}\alpha}({\color[rgb]{0,0,0}\bm{z}},{\color[rgb]{0,0,0}\bm{x}})=\exp\left[{\color[rgb]{0,0,0}k}^{\top}({\color[rgb]{0,0,0}\bm{z}}){\color[rgb]{0,0,0}q}({\color[rgb]{0,0,0}\bm{x}})\right]$,  where ${\color[rgb]{0,0,0}k}:\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{z}}}\rightarrow\mathbb{R}^{{\color[rgb]{0,0,0}D}}$ and ${\color[rgb]{0,0,0}q}:\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{x}}}\rightarrow\mathbb{R}^{{\color[rgb]{0,0,0}D}}$ are (typically linear) learnable functions that respectively return the so-called key and query vectors. One can think of ${\color[rgb]{0,0,0}k}({\color[rgb]{0,0,0}\bm{z}})$ as a vector that encodes the information in ${\color[rgb]{0,0,0}\bm{z}}$ and of ${\color[rgb]{0,0,0}q}({\color[rgb]{0,0,0}\bm{x}})$ as a vector that encodes the information relevant to ${\color[rgb]{0,0,0}\bm{x}}$. Thereby, ${\color[rgb]{0,0,0}\alpha}({\color[rgb]{0,0,0}\bm{z}},{\color[rgb]{0,0,0}\bm{x}})$ captures how relevant ${\color[rgb]{0,0,0}\bm{z}}$ is to ${\color[rgb]{0,0,0}\bm{x}}$.
@@ IV Attention Based RME
A simple possibility to design a transformer-based RME estimator would be to consider the existing DNN-based estimators and replace the CNN therein with a vision transformer [24]. However, this would suffer from the limitations of these estimators; cf. Sec. I. Instead, an alternative route is taken here, which proceeds by adopting a somehow abstract perspective. In particular, note from the problem formulation in Sec. II that any estimator of ${{\color[rgb]{0,0,0}\gamma}}({\color[rgb]{0,0,0}\bm{r}})$ given the data $\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}$ is a function of the form ${{\color[rgb]{0,0,0}\hat{{\color[rgb]{0,0,0}\gamma}}}}({\color[rgb]{0,0,0}\bm{r}};\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}})$.

### v2

@@ Abstract
Radio map estimation (RME) involves spatial interpolation of radio measurements to predict metrics such as the received signal strength at locations where no measurements were collected. The most popular estimators nowadays project the measurement locations onto a regular grid and complete the resulting measurement tensor with a convolutional deep neural network. Unfortunately, these approaches suffer from poor spatial resolution and require a very large number of parameters. The first contribution of this paper addresses these limitations by means of an attention-based estimator named Spatial TransfOrmer for Radio Map estimation (STORM). This scheme not only outperforms the existing estimators, but also exhibits lower computational complexity, translation equivariance, rotation equivariance, and full spatial resolution. The second contribution is an extended transformer architecture that allows STORM to perform active sensing, by which the next measurement location is selected based on the previous measurements. This is particularly useful for minimization of drive tests (MDT) in cellular networks, where operators request user equipment to collect measurements. Finally, STORM is extensively validated by experiments with one ray-tracing and two real-measurement datasets.
@@
Radio map estimation, transformers, attention-based learning, deep learning, wireless communications. This work has been funded by the IKTPLUSS grant 311994 of the Research Council of Norway.
@@ I Introduction
Radio maps (see Fig. 1), also known as radio environment maps, provide radio frequency (RF) metrics such as the received signal strength across a geographical region [1, 2]. Radio maps find a large number of applications, including network planning, frequency planning, cellular communications, device-to-device communications, dynamic spectrum access, robot path planning, aerial traffic management in unmanned aerial systems, and fingerprinting localization to name a few; see e.g. [3, 4, 1, 5] and references therein.
Radio map estimation (RME) involves constructing a radio map by relying on measurements collected across the area of interest. Before the advent of deep learning, the most popular estimators were built upon kernel-based learning (see [6] and references therein), Kriging [7, 8], sparsity-based inference [9], matrix completion [10], dictionary learning [11], and graphical models [12]. The most recent estimators are based on deep neural networks (DNNs); see e.g. [13, 14, 15, 8]. Unfortunately, these schemes entail grid discretization and a very large number of trainable parameters, which render them computationally expensive and drastically limit their spatial resolution. Besides, they lack important desirable properties in RME, such as translation and rotation equivariance.
@@
Some works related to RME have considered attention-based estimators, which is the topic of this paper. For example, vision transformers have been used to accommodate side information, such as building maps [16, 17] and satellite images [18]. In other works, transformers are fed with radio measurements. This is the case of [19], which uses a transformer for predicting what a device would measure given the measurements collected by a device with different hardware characteristics, and of [20], where a transformer fills missing RSS features for fingerprinting-based localization. Thus, transformers have been applied to problems that are related to RME or to modified versions of the RME problem where transformers are used to process images. However, the plain RME problem, where a radio map needs to be constructed by relying solely on radio measurements and their locations, requires spatial interpolation of radio measurements and this has never been tackled using transformers.
Figure 1: Example of a radio map estimate obtained with the proposed STORM estimator in the USRP dataset; cf. Sec. VI.
The first contribution of this paper is an attention-based scheme referred to as Spatial TransfOrmer for Radio Map estimation (STORM). As shown by numerical experiments in one ray-tracing and two real-measurement datasets, STORM outperforms the existing estimators, thereby setting the state of the art in RME. Besides, it offers key advantages over existing DNN estimators: First, its complexity is significantly lower. For example, the RadioUnet estimators in [14] have 6 M and 25 M parameters, whereas STORM has just 100 k parameters. Second, STORM is gridless. This implies that, unlike previous DNN estimators, (i) its spatial resolution is not limited, (ii) it is translation and rotation equivariant, which are desirable properties in view of Maxwell’s equations, and (iii), it need not be retrained e.g. if the grid spacing changes. Third, STORM can estimate the map only at the relevant locations, whereas previous DNN estimators need to compute the map at all grid locations. Fourth, STORM can accommodate measurements outside the region of interest, as opposed to existing DNN estimators. STROM also compares favorably to Kriging, which is the non-DNN estimator of choice [21, 22]. While Kriging requires a matrix inversion, whose complexity is cubic in the number of measurements, STORM enjoys quadratic complexity.
@@ II The RME Problem
This section presents the most prevalent formulation of RME, both in the conventional and active sensing setups. For simplicity, the signal strength is quantified here by the received signal power, but other metrics can readily be adopted.
@@
The received power is measured at ${{\color[rgb]{0,0,0}N}}$ locations $\left\{{\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}\subset{\color[rgb]{0,0,0}\mathcal{R}}$ by one or multiple receivers (or sensors). The ${{\color[rgb]{0,0,0}n}}$-th measurement can be written as ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}}={{\color[rgb]{0,0,0}\gamma}}\left({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\right)+{\color[rgb]{0,0,0}\zeta}_{{{\color[rgb]{0,0,0}n}}}$, where ${\color[rgb]{0,0,0}\zeta}_{{{\color[rgb]{0,0,0}n}}}$ denotes measurement error. Given $\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}$, the RME problem is to estimate ${{\color[rgb]{0,0,0}\gamma}}({\color[rgb]{0,0,0}\bm{r}})$, ${\color[rgb]{0,0,0}\bm{r}}\in{\color[rgb]{0,0,0}\mathcal{R}}$. On the other hand, in the active sensing problem, one is given the measurements $\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}$ as well as the set $\{{\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}}\}_{{{\color[rgb]{0,0,0}n}}={{\color[rgb]{0,0,0}N}}+1}^{{{\color[rgb]{0,0,0}N}}+{\color[rgb]{0,0,0}M}}$ of candidate locations and selects one of these candidate locations, say ${\color[rgb]{0,0,0}\bm{r}}_{{\color[rgb]{0,0,0}m}}$, to collect the next measurement ${\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{\color[rgb]{0,0,0}m}}$. The goal is to choose ${\color[rgb]{0,0,0}m}$ so that a target metric of the estimation error is minimized given the measurements $\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}\cup\{({\color[rgb]{0,0,0}\bm{r}}_{{\color[rgb]{0,0,0}m}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{\color[rgb]{0,0,0}m}})\}$.
@@ III-A Attention Heads
The attention weights are determined by the relation between ${\color[rgb]{0,0,0}\bm{x}}$ and the vectors in ${\color[rgb]{0,0,0}\bm{Z}}$. The learnable function ${\color[rgb]{0,0,0}\alpha}:\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{z}}}\times\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{x}}}\rightarrow\mathbb{R}_{++}$ can be thought of as quantifying the similarity between ${\color[rgb]{0,0,0}\bm{z}}$ and ${\color[rgb]{0,0,0}\bm{x}}$. In this way, if ${\color[rgb]{0,0,0}\bm{x}}$ is very similar to ${\color[rgb]{0,0,0}\bm{z}}_{{{\color[rgb]{0,0,0}n}}_{0}}$ for some ${{\color[rgb]{0,0,0}n}}_{0}$ and dissimilar to the remaining reference vectors, then ${\color[rgb]{0,0,0}\alpha}({\color[rgb]{0,0,0}\bm{z}}_{{{\color[rgb]{0,0,0}n}}_{0}},{\color[rgb]{0,0,0}\bm{x}})$ will dominate and ${\color[rgb]{0,0,0}H}({\color[rgb]{0,0,0}\bm{Z}},{\color[rgb]{0,0,0}\bm{x}})\approx{\color[rgb]{0,0,0}v}({\color[rgb]{0,0,0}\bm{z}}_{{{\color[rgb]{0,0,0}n}}_{0}})$. Usually, ${\color[rgb]{0,0,0}\alpha}$ is the so-called inner-product attention function [^1] A factor of $1/\sqrt{{\color[rgb]{0,0,0}D}}$ is often explicitly included inside the exponential but it is absorbed here into either ${\color[rgb]{0,0,0}k}$ or ${\color[rgb]{0,0,0}q}$ to simplify notation. ${\color[rgb]{0,0,0}\alpha}({\color[rgb]{0,0,0}\bm{z}},{\color[rgb]{0,0,0}\bm{x}})=\exp\left[{\color[rgb]{0,0,0}k}^{\top}({\color[rgb]{0,0,0}\bm{z}}){\color[rgb]{0,0,0}q}({\color[rgb]{0,0,0}\bm{x}})\right]$,  where ${\color[rgb]{0,0,0}k}:\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{z}}}\rightarrow\mathbb{R}^{{\color[rgb]{0,0,0}D}}$ and ${\color[rgb]{0,0,0}q}:\mathbb{R}^{{\color[rgb]{0,0,0}D}_{\bm{x}}}\rightarrow\mathbb{R}^{{\color[rgb]{0,0,0}D}}$ are (typically linear) learnable functions that respectively return the so-called key and query vectors. One can think of ${\color[rgb]{0,0,0}k}({\color[rgb]{0,0,0}\bm{z}})$ as a vector that encodes the information in ${\color[rgb]{0,0,0}\bm{z}}$ and of ${\color[rgb]{0,0,0}q}({\color[rgb]{0,0,0}\bm{x}})$ as a vector that encodes the information relevant to ${\color[rgb]{0,0,0}\bm{x}}$. Thereby, ${\color[rgb]{0,0,0}\alpha}({\color[rgb]{0,0,0}\bm{z}},{\color[rgb]{0,0,0}\bm{x}})$ captures how relevant ${\color[rgb]{0,0,0}\bm{z}}$ is to ${\color[rgb]{0,0,0}\bm{x}}$.
@@ IV Attention Based RME
A simple possibility to design a transformer-based RME estimator would be to consider the existing (grid-aware) DNN-based estimators and replace the DNN therein with a vision transformer [24]. However, this would suffer from the limitations of these estimators; cf. Sec. I. Instead, an alternative route is taken here, which proceeds by adopting a somehow abstract perspective. In particular, note from the problem formulation in Sec. II that any estimator of ${{\color[rgb]{0,0,0}\gamma}}({\color[rgb]{0,0,0}\bm{r}})$ given the data $\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}}$ is a function of the form ${{\color[rgb]{0,0,0}\hat{{\color[rgb]{0,0,0}\gamma}}}}({\color[rgb]{0,0,0}\bm{r}};\left\{({\color[rgb]{0,0,0}\bm{r}}_{{{\color[rgb]{0,0,0}n}}},{\color[rgb]{0,0,0}\tilde{{{\color[rgb]{0,0,0}\gamma}}}}_{{{\color[rgb]{0,0,0}n}}})\right\}_{{{\color[rgb]{0,0,0}n}}=1}^{{{\color[rgb]{0,0,0}N}}})$.